Remember to explain all your steps and to show all formulas you use before you put numbers into the equations.
-Write the formula, substitute the numbers into the formula, say what is left to compute, compute. Question 1 For question 1, use the following cash flows for projects A and B: A: (-$2,000, $500, $600, $700, $800) B: (-$2,
Answer #1

Answer to 1.1:

Information given in the question is cash flows:

Cash outflows

Project A = -$2,000 and project B = -$2,000

Cash inflows

Project A; $500, $600, $700, $800 in year 1, 2,3,4 respectively

Project B; $950, $850, $400, $300 in year 1,2,3,4 respectively

Payback period is period in which we can be able to recover our initial cash outflow/investment

Payback period for Project A:

Sum of year 1,2,3($500+$600+$700=$1800) cash inflows were lower than $2,000.

That means payback period is more than 3 years

Payback period = 3 + 200/800

= 3 + 0.25 = 3.25

Payback period for project A(cash flows -$2,000,$500, $600, $700, $800 ) is 3.25 years

Payback period for Project B:

Sum of year 1,2 ($950+$850 =$1800) cash inflows were lower than $2,000.

That means payback period is more than 2 years

Payback period = 2 + 200/400

= 2 + 0.5 = 2.5

Payback period for project B(cash flows -$2,000,$950, $850, $400, $300 ) is 2.5 years